Second International Workshop on Testing and Modeling (TEST 2003)

at the 22nd International Conference on Conceptual Modeling (ER 2003)

Chicago, Illinois, October 13, 2003

Organized by Peter A. Smith and Carl Miller

Preface to TEST2003

Springer asks that we write the preface in past tense since most readers will see it after the conference. The preface of your workshop should be between 1 and 2 pages.

– Example of some item list

– Feel free to have any number of items to explain the topics of interest

You should shortly introduce all papers presented at your workshop.

October 2003 Peter A. Smith Carl Miller